kayleeb:

How can you generate x + 3y + 5 from the expression x + y + 4 + 2y + 1? A. by using the Commutative Property B. by combining like terms C. by using the Associative Property D. by using the Distributive Property

lilianamendez2:

which one u think is correct

kayleeb:

c

lilianamendez2:

well c is wrong You need to combine like terms. you regroup all the x', y' and numbers together it will result in the expression x + 3y + 5. (x) = x (y) + (2y) = 3y (4) + (1) = 5. They are all positive, if u add them and your final simplify answer is x + 3y + 5 i hope it

lilianamendez2:

so ur answer is somewhere there